# LaserEscape: Detecting and Mitigating Optical Probing Attacks

### v1

@@ Abstract.
The security of integrated circuits (ICs) can be broken by sophisticated physical attacks relying on failure analysis methods. Optical probing is one of the most prominent examples of such attacks, which can be accomplished in a matter of days, even with limited knowledge of the IC under attack. Unfortunately, few countermeasures are proposed in the literature, and none has been fabricated and tested in practice. These countermeasures usually require changing the standard cell libraries and, thus, are incompatible with digital and programmable platforms, such as field programmable gate arrays (FPGAs). In this work, we shift our attention from preventing the attack to detecting and responding to it. We introduce LaserEscape, the first fully digital and FPGA-compatible countermeasure to detect and mitigate optical probing attacks. LaserEscape incorporates digital delay-based sensors to reliably detect the physical alteration on the fabric caused by laser beam irradiations in real time. Furthermore, as a response to the attack, LaserEscape deploys real-time hiding approaches using randomized hardware reconfigurability. It realizes 1) moving target defense (MTD) to physically move the sensitive circuity under attack out of the probing field of focus to protect secret keys and 2) polymorphism to logically obfuscate the functionality of the targeted circuit to counter function extraction and reverse engineering attempts. We demonstrate the effectiveness and resiliency of our approach by performing optical probing attacks on protected and unprotected designs on a 28-nm FPGA. Our results show that optical probing attacks can be reliably detected and mitigated without interrupting the chip’s operation.
@@ 1. Introduction
The security of a computing machine in various applications relies on sensitive assets stored on Root-of-Trust integrated circuits (ICs). Examples of on-chip assets include cryptographic keys, proprietary firmware, and intellectual property (IP). Such assets can be extracted by attackers, who can gain physical access to these devices and mount physical attacks. Attacks relying on failure analysis (FA) tools are the most threatening ones against ICs. The prominent examples of such attacks are the optical probing attack variants (Lohrke et al., 2016; Tajik et al., 2017b; Chef et al., 2018; Krachenfels et al., 2021a, b) which enable the adversary to recover secret data and functions from chips. Illuminating a laser beam on transistors and measuring their states using the modulated backscattered laser beam is the main mechanism behind optical probing attacks.
@@
Deploying randomization in countermeasures, e.g., as masking (Nikova et al., 2006; Groß et al., 2016), is a conventional method to mitigate conventional side-channel attacks, as it prevents the repetition and integration of the measurements over multiple clock cycles. However, randomness becomes ineffective if the adversary performs optical probing on halted circuits or data at rest (Krachenfels et al., 2021a, b). Moreover, many root/master keys, which cannot be masked, are vulnerable to optical probing during key caching from memory to registers (Lohrke et al., 2016; Krachenfels et al., 2021b).
@@
A few prevention and detection-based countermeasures have been proposed in the literature to avert optical probing attacks. These countermeasures try to protect the design either directly from the laser irradiation (Rahman et al., 2021; Parvin et al., 2023; Zhang et al., 2023) or detect tampering with voltage/clock, which is the required step for some specific class of optical probing attacks (Farheen et al., 2022). Such schemes deploy various logic styles and design techniques to either minimize the data/function-dependent leakage through the backscattered laser light (Rahman et al., 2021; Parvin et al., 2023) or to detect the generated heat illumination of the laser beam on logic cells (Zhang et al., 2023). Most of these solutions require the modification of standard cell libraries and are incompatible with legacy systems and field programmable gate arrays (FPGAs). Moreover, they are tailored toward a specific class of optical probing and, thus, cannot mitigate the threat in general. Finally, these schemes have yet to be fabricated or tested in practice.
@@
The only FPGA-compatible sensing countermeasure that has been proposed so far deploys the ring-oscillators (ROs) (Tajik et al., 2017a) to detect laser irradiation. However, such a solution has never been researched further due its high sensitivity to the environmental conditions, which leads to high false alarms. In this work, we attempt to answer the following research questions: (1) Is it possible to confidently detect the optical probing attempt in real-time using fully digital components of an FPGA? (2) What would be a suitable response upon detecting an optical probing attempt to deceive the adversary and not interrupt the IC’s operation?
@@
Our Contribution. To answer the above questions, we present LaserEscape, which is capable of detecting and responding to optical probing attacks using the existing resources of mainstream FPGAs. To detect an optical probing attempt, we repurpose the recently proposed fully digital delay-based 1LUTSensor (Jayasinghe et al., 2023), which only needs one single look-up table (LUT) to detect the laser irradiation from the IC backside silicon. Afterward, to respond to the attack, we explore novel moving target defense and gate polymorphism strategies to deceive the adversary in real-time and render the attack ineffective. First, we demonstrate that how upon the detection of the attack, the partial reconfiguration (PR) feature of mainstream FPGAs and programmable SoCs can be utilized to dynamically randomize the placement of the security-critical registers and moving them out from attacker’s probing field with a low overhead to protect secret keys. Second, we show how polymorphic logic gates can be realized on LUTs of FPGAs to zeroize the output of the logic function to prevent function extraction by the attacker. Finally, we apply LaserEscape to various circuits on a 28-nm FPGA and perform optical probing on them to assess the effectiveness of our solution.
@@ 2.1. Optical Probing Attacks
Fig. 1 illustrates the high-level overview of optical probing attacks. As illustrated, the laser scanning microscope can be used to navigate through a target IC as well as capture dynamic transistor-level activities on the chip. Optical probing attacks often include extracting sensitive data from the target chip while IC is operational (Lohrke et al., 2016; Tajik et al., 2017b; Chef et al., 2018; Krachenfels et al., 2021b, a). For this aim, a high-power incoherent light source (HIL) with the wavelength of 1.3 $\mu$m is utilized. In 1, the light beams are illuminated and passed through multiple optical mirrors, including the scanning mirrors, before entering the objective lenses. The illumined beams are reflected from the IC, and are guided to the laser detector in 2 using splitters. By focusing the objective lens onto the region of interest (RoI), 2D image of the target area can be captured and then be used for localizing the physical components on the IC. Additionally by moving the stage different parts of the IC can be navigated.
@@ 2.2. Sensor Implementations on FPGAs
In the case of laser attacks, such as optical probing, the colliding laser photons locally increase the temperature of the target wires and transistors on the chip. The local temperature variations can affect the propagation delays of the electrical signals in the delay-based sensors on FPGAs, such as ring-oscillators and time-to-digital converters (TDCs). ROs (Zhao and Suh, 2018) and TDCs (Schellenberg et al., 2021) have been extensively used for remote side-channel attacks on FPGAs, where they measure the voltage fluctuations on the power delivery network (PDN) caused by adjacent intellectual property cores. While these sensors are often used for offensive side-channels, they can also be used for the detection of active attacks causing temperature and current variations. In an attempt, RO networks have been employed to detect optical probing attacks on FPGAs (Tajik et al., 2017a). The main drawbacks of using RO networks are their high power consumption and sensitivity to environmental conditions, leading to a high number of false alarms(Tajik et al., 2017a). Therefore, as described later, a more reliable TDC candidate will be chosen to detect optical probing attacks.
@@ 2.3. Placement Randomization as MTD
Several countermeasures against side-channel attacks, like power and electromagnetic (EM) analysis, have already leveraged the PR (Mentens et al., 2008; Güneysu and Moradi, 2011; Heyszl et al., 2012; Moradi and Mischke, 2013; Hettwer et al., 2019; Bow et al., 2020; Khan et al., 2021). Some of these approaches use PR to introduce timing jitter, which can help distort power traces and defeat power side-channel attacks. Other techniques involve dynamically relocating functional units to different regions of the FPGA through PR, dissociating the spatial correlation between EM emanations and the target logic. The capability for real-time reconfiguration of FPGAs through PR has also been explored to create side-channel resistant implementations for cryptographic primitives, such as (Monfared et al., 2024).
@@ 3. Threat Model and Attack Scenarios
For optical probing attacks, it is presumed that an adversary is equipped with FA tools and aims to extract a secret key or element-specific reverse engineering. The primary motivation often are in the scenario where a single key is employed across all devices, such as in instances where firmware, bitstream, or logic encryption secures the proprietary design of a system. The vendor programs this key into the product before distribution to customers. Thus, acquiring the key from one device could compromise the security of the entire device family. Moreover, even if individual devices use different keys, the uniform layout across a device family means that once an adversary masters key extraction from one device, she can apply this knowledge to breach other devices within the same family. At the right side of the Fig. 1, attack scenarios are highlighted in 7. For this article we consider optical-based 1) key recovery attacks(Krachenfels et al., 2021b, a), and 2) function recovery attempts(Mehta et al., 2024).
@@ 4.2. Sensor Architecture
The primary reason behind choosing this sensor is its size compared to other TDCs, which only require one LUT and can be placed in any slice, satisfying the spatial coverage requirements. Furthermore, the utilization of these PVT-invariant elements for delay tuning makes the sensor immune to environmental conditions, satisfying its reliability. Finally, supplying the sensor a few tens of megahertz clock signal provides sufficient timing resolution less than the laser dwell time on a pixel.
@@ 4.3. Adjustable Delay Elements
IDELAYE2 elements are 31-tap, wraparound, delay element with a calibrated tap resolution (Xilinx, 2023). Each IDELAYE2 acts as a pass-through for a 1-bit signal with a 5-bit input for the delay configuration, where the higher value leads to a larger delay. Since the delay of the clock path is larger than that of the data path, a single IDELAYE2 might not provide sufficient delay. Therefore, a chain of IDELAYE2s for the clock path can be used as shown in Fig. 4.
@@ 4.4. Sensor Tuning
In our sensor implementation, we use a standalone IDELAYE2 for the data path delay and an $m-$long IDELAYE2 chain for the clock path. Moreover, we use a select input to determine which input of the LUT goes through to the register. Hence, we establish a tune value, consisting of three values: 1) data-path delay, 2) clock-path delay, and 3) select. For many combinations of data path delay and LUT select, there is a corresponding value for the clock path delay where the output is metastable. However, some combinations perform better since their accumulative zero counts are significantly smaller than others. For every value of the data path delay, the binary search of clock delay for metastability is carried out with LUT select fixed at zero. If metastability is found, we search for adjacent clock delay values with every LUT select value for other metastable tunes. During this process, we keep track of the best tune, determined by the lowest maximum zero count over the $t_{sense}$ interval. At the end of the procedure, we extract the best tune.
Figure 4. Block diagram of a 5-long IDELAYE2 chain
@@ 4.5. Sensor Output Post Processing
When the sensor’s output is metastable, yielding an output of $1$ with occasional zero pulses. The length and frequency of these pulses increase during optical probing due to the impact of increased temperature on delay lines. Hence, we utilize measurement circuitry to characterize this. First, we implement a zero counter, which counts the number of clock cycles for which the sensor value is zero over $t_{detect}$. Secondly, we employ a pulse counter, which counts the length of the zero pulses. Based on the readings of the counters, we connect the sensor output to a latch, which we use to trigger our chosen defenses: 1) partial reconfiguration or 2) polymorphic gates, which will be described in Sect 5.1 and Sect 5.2, respectively.
@@ 5. Attack Response
Real-time response against physical attacks can addressed with different approaches in accordance to the threat model. For instance one can deploy a self-destructive mechanism to zeroize the sensitive data upon sensing the physical attacks (Cannon et al., 2023; Tada et al., 2021). In contrast to solutions which necessitates cell library modification at pre-silicon (Shen et al., 2018), we propose fully logic-level mitigation methodologies that provides non-interrupting circumvention in addition to existing CMOS friendly zeroization (Tada et al., 2021; Cannon et al., 2023). Specifically, in our method target circuitry can benefit from real-time reconfiguration to bypass the optical attack and almost uninterrupted, operate as normal. This is useful for deceiving the adversary (Tajik et al., 2017a). Furthermore, the full zeroization of the device including the entire memory contents incurs a huge area and latency for large designs, (Cannon et al., 2023; Srivastava and Ghosh, 2019) which can be replaced by partial zeroization. In addition to MTD, we incorporate a gate polymorphic design, particularly to defeat optical reverse engineering attempts. We design the polymorphic gates by inspiring the state-of-the-art methodology in (Cannon et al., 2023) to create a supply voltage-controlled gates and obfuscate the functionality of the target elements in real-time.

### v2

@@ Abstract.
The security of integrated circuits (ICs) can be broken by sophisticated physical attacks relying on failure analysis methods. Optical probing is one of the most prominent examples of such attacks, which can be accomplished in a matter of days, even with limited knowledge of the IC under attack. Unfortunately, few countermeasures are proposed in the literature, and none have been fabricated and tested in practice. These countermeasures usually require changing the standard cell libraries and, thus, are incompatible with digital and programmable platforms, such as field programmable gate arrays (FPGAs). In this work, we shift our attention from preventing the attack to detecting and responding to it. We introduce LaserEscape, the first fully digital and FPGA-compatible countermeasure to detect and mitigate optical probing attacks. LaserEscape incorporates digital delay-based sensors to reliably detect the physical alteration of the fabric caused by laser beam irradiations in real time. Furthermore, as a response to the attack, LaserEscape deploys real-time hiding approaches using randomized hardware reconfigurability. It realizes 1) moving target defense (MTD) to physically move the sensitive circuity under attack out of the probing field of focus to protect secret keys and 2) polymorphism to logically obfuscate the functionality of the targeted circuit to counter function extraction and reverse engineering attempts. We demonstrate the effectiveness and resiliency of our approach by performing optical probing attacks on protected and unprotected designs on a 28-nm FPGA. Our results show that optical probing attacks can be reliably detected and mitigated without interrupting the chip’s operation.
@@
journalyear: 2024copyright: acmlicensedconference: IEEE/ACM International Conference on Computer-Aided Design; October 27–31, 2024; New York, NY, USAbooktitle: IEEE/ACM International Conference on Computer-Aided Design (ICCAD ’24), October 27–31, 2024, New York, NY, USAdoi: 10.1145/3676536.3676822isbn: 979-8-4007-1077-3/24/10
@@ 1. Introduction
In various applications, the security of a computing machine relies on sensitive assets stored on Root-of-Trust integrated circuits (ICs). Examples of these on-chip assets include cryptographic keys, proprietary firmware, and intellectual property (IP). Attackers can extract these assets by gaining physical access to the devices where they are stored and mounting physical attacks. Attacks relying on failure analysis (FA) tools pose the biggest threat against ICs. Prominent examples of such attacks are the various forms of optical probing attacks (Lohrke et al., 2016; Tajik et al., 2017b; Chef et al., 2018; Krachenfels et al., 2021a, b) which enable the adversary to recover secret data and functions from chips. Illuminating a laser beam on transistors and measuring their states using the modulated backscattered laser beam is the main mechanism behind these optical probing attacks.
@@
Deploying randomization in countermeasures, such as masking (Nikova et al., 2006; Groß et al., 2016), is a conventional method to mitigate conventional side-channel attacks, as it prevents the repetition and integration of the measurements over multiple clock cycles. However, randomness becomes ineffective if the adversary performs optical probing on halted circuits or data at rest (Krachenfels et al., 2021a, b). Moreover, many root/master keys, which cannot be masked, are vulnerable to optical probing during key caching from memory to registers (Lohrke et al., 2016; Krachenfels et al., 2021b).
@@
A few prevention and detection-based countermeasures have been proposed in the literature to avert optical probing attacks. These countermeasures try to protect the design either directly from laser irradiation (Rahman et al., 2021; Parvin et al., 2023; Zhang et al., 2023) or detect tampering with the voltage/clock, which is required for certain optical probing attacks (Farheen et al., 2022). Such schemes deploy various logic styles and design techniques to either minimize the data/function-dependent leakage through the backscattered laser light (Rahman et al., 2021; Parvin et al., 2023) or to detect the generated heat illumination of the laser beam on logic cells (Zhang et al., 2023). Most of these solutions require the modification of standard cell libraries and are incompatible with legacy systems and field programmable gate arrays (FPGAs). Moreover, they are tailored toward a specific form of optical probing and, thus, cannot mitigate the threat in general. Finally, these schemes have yet to be fabricated or tested in practice.
@@
The only FPGA-compatible sensing countermeasure that has been proposed so far deploys ring-oscillators (ROs) (Tajik et al., 2017a) to detect laser irradiation. However, such a solution has never been researched further due its high sensitivity to environmental conditions, which leads to high false alarms. In this work, we attempt to answer the following research questions: (1) Is it possible to confidently detect optical probing attempts in real-time using fully digital components of an FPGA? (2) What would be a suitable response upon detecting an optical probing attempt to deceive the adversary and not interrupt the IC’s operation?
@@
Our Contribution. To answer the above questions, we present LaserEscape, which is capable of detecting and responding to optical probing attacks using the existing resources of mainstream FPGAs. To detect an optical probing attempt, we repurpose the recently proposed fully digital delay-based 1LUTSensor (Jayasinghe et al., 2023), which had been originally been designed for remote power analysis side-channel attacks. We found that the mechanism which allows 1LUTSensor to detect power fluctuations with only one look-up table (LUT) can also detect laser irradiation from the IC backside silicon. Afterward, to respond to the attack, we explore novel moving target defense and gate polymorphism strategies to deceive the adversary in real-time and render the attack ineffective. First, we demonstrate that how upon the detection of the attack, the partial reconfiguration (PR) feature of mainstream FPGAs and programmable SoCs can be utilized to dynamically randomize the placement of security-critical registers and moving them out from attacker’s probing field with a low overhead to protect secret keys. Second, we show how polymorphic logic gates can be realized on LUTs of FPGAs to zeroize the output of the logic function to prevent function extraction by the attacker. Finally, we apply LaserEscape to various circuits on a 28-nm FPGA and perform optical probing on them to assess the effectiveness of our solution.
Source Code Availability. We publicly publish our raw sensor data and make the source code of LaserEscape open source in https://github.com/vernamlab/LaserEscape
@@ 2.1. Optical Probing Attacks
Fig. 1 illustrates a high-level overview of optical probing attacks. As illustrated, a laser scanning microscope can be used to navigate through a target IC as well as capture dynamic transistor-level activities on it. Optical probing attacks often include extracting sensitive data from the target IC while it is operational (Lohrke et al., 2016; Tajik et al., 2017b; Chef et al., 2018; Krachenfels et al., 2021b, a). For this aim, a high-power incoherent light source (HIL) with a wavelength of 1.3 $\mu$m is utilized. In 1, the light beams are illuminated and passed through multiple optical mirrors, including scanning mirrors, before entering the objective lenses. The illumined beams are reflected from the IC, and are guided to the laser detector in 2 using splitters. By focusing the objective lens onto the region of interest (RoI), a 2D image of the target area can be captured and then be used for localizing physical components on the IC. Additionally by moving the stage different parts of the IC can be navigated.
@@ 2.2. Sensor Implementations on FPGAs
In the case of laser attacks, such as optical probing, the colliding laser photons locally increase the temperature of the target wires and transistors on the chip. These local temperature variations can affect the propagation delays of electrical signals in delay-based sensors on FPGAs, such as ring-oscillators and time-to-digital converters (TDCs). ROs (Zhao and Suh, 2018) and TDCs (Schellenberg et al., 2021) have been extensively used for remote side-channel attacks on FPGAs, where they measure voltage fluctuations on the power delivery network (PDN) caused by adjacent intellectual property cores. While these sensors are often used for offensive side-channels, they can also be used for the detection of active attacks causing temperature and current variations. RO networks have been employed as the first attempt to detect optical probing attacks on FPGAs (Tajik et al., 2017a). The main drawbacks of using RO networks are their high power consumption and sensitivity to environmental conditions, leading to a high number of false alarms(Tajik et al., 2017a). Therefore, as described later, a more reliable TDC candidate will be chosen to detect optical probing attacks.
@@ 2.3. Placement Randomization as MTD
Several countermeasures against side-channel attacks, like power and electromagnetic (EM) analysis, have already leveraged PR (Mentens et al., 2008; Güneysu and Moradi, 2011; Heyszl et al., 2012; Moradi and Mischke, 2013; Hettwer et al., 2019; Bow et al., 2020; Khan et al., 2021). Some of these approaches use PR to introduce timing jitter, which can help distort power traces and defeat power side-channel attacks. Other techniques involve dynamically relocating functional units to different regions of the FPGA through PR, dissociating the spatial correlation between EM emanations and the target logic. The capability for real-time reconfiguration of FPGAs through PR has also been explored to create side-channel resistant implementations for cryptographic primitives, such as (Monfared et al., 2024).
@@ 3. Threat Model and Attack Scenarios
For optical probing attacks, we presume that an adversary is equipped with FA tools and aims to extract a secret key or reverse engineer specific elements. The primary motivation for these attacks comes from the scenario where a single key is employed across all devices, such as in instances where firmware, bitstream, or logic encryption secures the proprietary design of a system. The vendor programs this key into the product before distribution to customers. Thus, acquiring the key from one device could compromise the security of the entire device family. Moreover, even if individual devices use different keys, the uniform layout across a device family means that once an adversary masters key extraction from one device, she can apply this knowledge to breach other devices within the same family. At the right side of the Fig. 1, attack scenarios are highlighted in 7. For this article we consider optical-based 1) key recovery attacks(Krachenfels et al., 2021b, a), and 2) function recovery attempts(Mehta et al., 2024).
@@ 4.2. Sensor Architecture
The primary reason behind choosing this sensor is its size compared to other TDCs, in that it only requires one LUT and can be placed in any slice, satisfying the spatial coverage requirements. Furthermore, the utilization of these PVT-invariant elements for delay tuning makes the sensor immune to environmental conditions, satisfying its reliability. Finally, supplying the sensor a few tens of megahertz clock signal provides sufficient timing resolution less than the laser dwell time on a pixel.
A key difference in our implementation is using a much lower clock frequency. The original 1LUTSensor as described in (Jayasinghe et al., 2023) was developed to measure voltage fluctuations in a circuit running on the chip to perform power analysis attacks. As such, it ran as fast as it could on the FPGA fabric - 600MHz - to measure an AES circuit running at frequencies ranging from 12-120MHz with enough resolution to detect the effects of switching gates. We do not want such a precise resolution, since this would cause the circuit we wish to protect to add noise to our readings. In addition, the laser pulses for LVI are on the order of milliseconds, which means the fluctuations in the sensor may be of that size.
@@ 4.3. Adjustable Delay Elements
IDELAYE2 elements are 31-tap, wraparound, delay elements with a calibrated tap resolution (Xilinx, 2023). Each IDELAYE2 acts as a pass-through for a 1-bit signal with a 5-bit input for the delay configuration, where the higher value leads to a larger delay. Since the delay of the clock path is larger than that of the data path, a single IDELAYE2 might not provide sufficient delay. Therefore, a chain of IDELAYE2s for the clock path can be used as shown in Fig. 4.
@@ 4.4. Sensor Tuning
In our sensor implementation, we use a standalone IDELAYE2 for the data path delay and an $m$-long IDELAYE2 chain for the clock path. Moreover, we use a select input to determine which input of the LUT goes through to the register. Hence, we establish a tune value, consisting of three values: 1) data-path delay, 2) clock-path delay, and 3) LUT select. For many combinations of data path delay and LUT select, there is a corresponding value for the clock path delay where the output is metastable. However, some combinations perform better since their accumulative zero counts are significantly smaller than others, which means they are metastable in a way that is closer to the constant zero case shown in Fig. 3(a). For every value of the data path delay, the binary search of clock delay for metastability is carried out with LUT select fixed at zero. If metastability is found, we search for adjacent clock delay values with every LUT select value for other metastable tunes. During this process, we keep track of the best tune, determined by the lowest maximum zero count over the $t_{sense}$ interval. At the end of the procedure, we extract the best tune.
Figure 4. Block diagram of a 4-long IDELAYE2 chain
@@ 4.5. Sensor Output Post Processing
When the sensor’s output is metastable, yielding an output of $1$ with occasional zero pulses, the length and frequency of these pulses increase during optical probing due to the impact of increased temperature on delay lines. Hence, we utilize measurement circuitry to characterize this. First, we implement a zero counter, which counts the number of clock cycles for which the sensor value is zero over $t_{detect}$. Secondly, we employ a pulse counter, which counts the length of the zero pulses. Based on the readings of the counters, we connect the sensor output to a latch, which we use to trigger our chosen defenses: 1) partial reconfiguration or 2) polymorphic gates, which will be described in Sect 5.1 and Sect 5.2, respectively.
@@ 5. Attack Response
Real-time response against physical attacks can be addressed with different approaches in accordance to the threat model. For instance one can deploy a self-destructive mechanism to zeroize the sensitive data upon sensing the physical attacks (Cannon et al., 2023; Tada et al., 2021). In contrast to solutions which necessitates cell library modification at pre-silicon (Shen et al., 2018), we propose fully logic-level mitigation methodologies that provides non-interrupting circumvention in addition to existing CMOS friendly zeroization (Tada et al., 2021; Cannon et al., 2023). Specifically, in our method target circuitry can benefit from real-time reconfiguration to bypass the optical attack and almost uninterrupted, operate as normal. This is useful for deceiving the adversary (Tajik et al., 2017a). Furthermore, the full zeroization of the device including the entire memory contents incurs a huge area and latency for large designs, (Cannon et al., 2023; Srivastava and Ghosh, 2019) which can be replaced by partial zeroization. In addition to MTD, we incorporate a gate polymorphic design, particularly to defeat optical reverse engineering attempts. We design the polymorphic gates by inspiring the state-of-the-art methodology in (Cannon et al., 2023) to create a supply voltage-controlled gates and obfuscate the functionality of the target elements in real-time.